A crew rowing with the current rowed 11 mi in 2 hours. Against the current, the crew rowed 5 mi in 2 hours. Find the rate of the current.
Answer by flame8855(424)

let x be the original speed of the crew
let c be the speed of the current
speed = distance/time
with hte current : speed = x +c
x+c = 11/2
with speed against current :
speed = x-c
x-c = 5/2
sove the two equations:
x = 4
c = 3/2 = 1.5 mph